# Enhancing Speech Large Language Models with Prompt-Aware Mixture of Audio Encoders

### v2

@@ 5 Discussions
Ablation study and routing method: To further validate the effectiveness of PaM, we conducted an ablation study, like PaM without the shared expert, as detailed in Appendix A.8. We also compare PaM against a learnable routing approach without task information, commonly employed in MoE models. The results indicate that PaM is more efficient and better suited to handling multiple downstream tasks than conventional routing and fusion strategies. Moreover, Table 3 highlights the strengths of each encoder. These findings encourage researchers to select encoders tailored to their downstream tasks for the Audio-LLM. While we did not incorporate this prior knowledge into the routing mechanism, leveraging it presents a promising direction for future work.
Other further work: 1.More and Unseen Tasks. Although our experiments involve three tasks and five datasets, they are representative as they encompass both semantic and acoustic-related tasks. We believe PaM can be extended to other tasks, which we will validate in future work. 2.Leverage Multimodal LLMs. Features across various encoders hinder initialization from pretrained multimodal LLMs in our work (Appendix A.9). We will explore strategies to more effectively leverage pretrained multimodal LLMs, such as Lai et al. (2024). 3.Efficient. Improving the efficiency of LLMs has attracted much attention. Enhancing the computational efficiency of speech LLMs presents another promising avenue for exploration (Appendix A.9).
@@ 6 Related Works
Audio encoders can be classified into supervised and self-supervised models. Supervised models typically employ ASR tasks to train an end-to-end model with an audio encoder and a text decoder. By omitting the decoder, the encoder can serve as a feature extractor Radford et al. (2022); Baevski et al. (2020). Self-supervised models can be trained on unlabeled speech signals. For instance, Wav2Vec2 Baevski et al. (2020) and HuBERT Hsu et al. (2021) were trained to predict the pseudo-discrete target at masked time steps. WavLM Chen et al. (2022) is a variant of HuBERT, designed to facilitate speaker identity extraction by using multi-speaker signals. Different model architectures, training methods, and data can result in encoders with distinct properties and advantages, making the mixture of audio encoders effective for Speech LLMs.
@@
To construct end-to-end speech LLMs, a natural approach is to extract discrete tokens from continuous speech signals and then expand the vocabulary of text LLMs to understand these speech tokens Rubenstein et al. (2023b); Veluri et al. (2024); Ma et al. (2024a). An alternative is to use an adapter layer to directly convert the continuous speech features into the continuous embedding space of the LLM. For example, QwenAudio Chu et al. (2024) employs average pooling to downsample speech features, followed by two linear layers for projection. SALMONN Tang et al. (2024) utilizes the Q-former Yu et al. (2024), a cross-attention-based adapter, to achieve a higher compression ratio. In parallel, to achieve high compression, Soundwave replaces cross-attention with a lightweight self-attention module that treats the inputs as query Zhang et al. (2025). Compared to previous works, our adapter handles more encoders and generates different features based on the prompt, rather than a single feature for all prompts.
@@
MoE has attracted growing interest, which replaces the FFN sub-layer in Transformer models with multiple experts Shazeer et al. (2017). These MoE methods typically employ massive experts and extremely sparse activation routing, increasing model size while maintaining constant inference costs, without explicitly considering the specialization of individual experts Fedus et al. (2022); Lepikhin et al. (2020). However, the vast scale of these models presents significant challenges for deployment. In contrast, the earliest MoE research introduced a data-dependent, trainable combining method Jacobs et al. (1991); Masoudnia and Ebrahimpour (2014), which aims to decompose complex tasks into simpler sub-tasks, each managed by a dedicated expert. Such works have inspired recent advances in developing modular models called expert specialization Ma et al. (2018); Gupta et al. (2022), providing solutions for deploying large-scale MoE models Lu et al. (2024) and enabling individual experts to learn and decompose diverse knowledge Dai et al. (2024). Inspired by these insights, we proposed a specialized MoE fusion method integrating multiple audio features to enhance Speech LLMs.
@@ 7 Conclusion
In conclusion, we propose PaM, a feature fusion method designed to provide Speech LLM with diverse features from multiple encoders based on users’ input prompts. Experimental results indicate that PaM surpasses both single-encoder and multi-encoder baselines across a variety of tasks and datasets. We provide a detailed analysis of the feature importance of different encoders, demonstrating that PaM effectively leverages different encoders and levels of features for distinct tasks. Additionally, we present comprehensive experimental results for the selection and combination of encoders. For future work, we intend to expand the training data and incorporate additional tasks.
@@ 8 Limitations
Owing to resource constraints, our training data is limited to several hundred hours. It would be preferable to implement our method in larger-scale experiments to facilitate comparison with existing strong Speech LLMs such as Qwen-Audio Chu et al. (2024) on a more comprehensive benchmark like AirBench Yang et al. (2024). Additionally, we train the PaM module from scratch using a predefined list of audio encoders. It would be beneficial to investigate the addition of new encoders to an already trained Speech LLM to enhance its performance on new tasks or in new domains. We leave this for future work.
@@ A.2 Details of Models and Datasets
In this paper, we leverage multiple audio encoders and LLM to construct the end-to-end speech LLM. Our training dataset is sourced from commonly used open-source datasets, totalling approximately 450 hours of audio data, corresponding to 313,208 samples, as outlined in Table 5. For SNV, we randomly concatenate individual utterances to form new speech signals with the number of speakers ranging from one to four.
@@
Table 5: The whole training dataset.
@@
A.3 Details of Baseline Implement
@@
For the single encoder baseline, we train the model using the same settings as in our method. For the second baseline, we train the model using the open-source codebases from WavLLM and SALMONN. We integrate the adapter components from these repositories into our code and train the baseline model using our training data, employing the same pre-trained audio encoders and LLMs as in our method. During training, we applied the same hyperparameters as our method. Since we use different encoders and LLMs compared to the baselines, we adjust the dimensions of the adapter to match the specific audio encoder and LLM we used while maintaining other dimensions independent of the audio encoders and LLM unchanged. Notably, we trained the SALMONN with query length=32 (as training with the original setting query length=1 failed) to ensure comparable performance with the other baseline methods.
@@ A.4 Results on Audio Caption with More Metrics
Due to the multiple evaluation metrics for the AC task, we scored the AC and AC QA tasks using more metrics in Table 7, including CIDEr, SPICE (with coco-caption toolkit), FENSE metrics, and Sentence-BERT [^8] We used the FENSE open-source repository GitHub and scored the entire dataset with eval_system.py. For the SBERT model, we loaded the paraphrase-mpnet-base-v2 model, and for the echecker, we used echecker_clotho_audiocaps_base. However, we encountered a bug when loading the echecker model, which had an unexpected key encoder.embeddings.position_ids. To resolve this, we set strict=False. Additionally, we included results based on similarity using the SBERT model.. We found that the different models exhibited similar performance trends across almost all metrics.
@@ A.5 PaM with More Encoders
We added the BEATs encoder to our framework (which includes four encoders) and found that it significantly improves the performance of our system on the AC task (Table 8). Although fusing the new encoder had some effect on the AMI and SNV tasks, incorporating the BEATs encoder improved the system’s average rank across downstream tasks (Table 9). We plan to conduct additional experiments with the EAT encoder Chen et al. (2024a) in further work.
@@ A.6 Details of Training and Inference Parameters
We train our model for five epochs with a learning rate of 5e-5, 2000 warmup steps, and bf16 precision. We freeze all encoders and the LLM, only training adapters and the fusion modules. For the LLM, we apply LoRA Hu et al. (2022) with a rank of 32 and an alpha of 64, adding LoRA only on the q_proj and k_proj. Each task has the same probability during training. During the inference stage, we select the last checkpoint on the validation set and perform greedy search.
@@ A.7 WavLLM and PaM with Larger LLM
Table 10: Results based on our PaM adapter and WavLLM with larger LLM.
Notably, we note that the concatenation method in the Llama-based model performs significantly worse on SNV, with only 3.1% accuracy. This is because it becomes difficult to follow the instructions of the SNV task during the inference stage. After further experiments, we found that the concatenation method becomes progressively less effective on SNV. This suggests that the method struggles to achieve a balance between multitasking as training progresses.
@@ A.8 Ablation Experiments of Routing Method
  • Audio-based Routing Method. The routing method employed in most of the MoE models, such as Switch Transfomers Fedus et al. (2022), DeepSeekMoE Dai et al. (2024), and MoWE Zhang et al. (2024), which use the current layer input as the routing module input, and optimize routing module directly based on the loss of outputs. However, ignores information from task labels.
@@ G $=\text{Top-}k(\text{Softmax}(\mathbf{X}))$ (11)
MoE (audio-based) 4.12 8.32 11.21 26.00% 35.66 15.61 3.17
PaM (with one expert) 4.27 10.04 13.77 45.80% 35.76 16.47 2.83
PaM (without shared) 4.31 7.89 13.43 45.20% 35.27 16.35 3.33
PaM (without task label) 3.97 7.97 13.14 43.50% 35.54 15.42 3.17
PaM (ours) 3.65 7.07 12.79 42.80% 35.47 15.70 2.50
Table 11: Ablation results on our routing method and the result based on the audio-based routing method
@@ A.9 Initialization setup for LLM
We train the LLM model from the open-source base model, following the setup used widely in prior work Hu et al. (2024); Ma et al. (2024b); Tang et al. (2024); Rubenstein et al. (2023a). An interesting alternative is to initialize our LLM module with the LLM module in a pretrained multimodal model like Qwen-audio. While we are not adapting such a setup because Qwen-Audio’s feature alignment was specifically designed for Whisper’s encoder, which can potentially "overfit" to features from Whisper. Our new experiments in Table 12 reveal that the feature spaces of Wav2Vec2 and WavLM are relatively similar, while Whisper’s feature space shows greater divergence. This pattern is also reflected in the weight distribution in Figure 3, where Wav2Vec2 and WavLM appear more closely aligned and significantly different from Whisper. Therefore, using other encoders and restructuring the projector module would still require re-adapting the LLM to comprehend new features.
@@ A.10 All Detailed Results
The detailed results of our experiments with multiple encoders are summarized in Table 13. We observe that, in most cases, the audio encoder that performs well on a single task also enhances the performance of the fusion model on that task. In cases where performance degradation occurs on a specific task when using the corresponding encoder, the fusion model consistently includes the HuBERT audio encoder, suggesting that incorporating the HuBERT model may have a detrimental effect. This could be attributed to the fact that the HuBERT model is trained on a smaller pre-trained dataset compared to other audio encoders. Notably, even in this case, fusing four audio encoders yields comparable results to fusing three encoders on the AVG Rank, indicating that incorporating more encoders can still lead to performance improvements.
@@ A.11 Prompts in Training and Inference Stage
In practice, large speech-language models typically address downstream tasks using prompts that are semantically explicit but textually diverse, as illustrated in Table 14. Unlike classical MoE routing methods that rely on hidden states, or other approaches such as predefined task labels (e.g., “Task ID = 3”) Kudugunta et al. (2021) or random routing Zuo et al. (2021), our model leverages natural language prompts to route inputs to the appropriate expert.
These prompts vary in phrasing but convey the same task intent. We utilize the semantic understanding capabilities of a pretrained LLM to extract task information from these prompts, rather than depending on fixed task labels. As an example, in our ASR task, we trained the routing module using 50 diverse prompts (Table 15). During evaluation, the model was tested on 200 prompts, 150 of which were unseen during training (Table 16). Our prompt-based router, powered by the LLM, achieved 100% expert selection accuracy, demonstrating strong generalization to previously unseen prompts.
Convert the audio speech into a text transcript.
Write an accurate version of the audio content in text.
Capture what is being said in this audio as text.
Table 14: Example of common prompts used in ASR tasks.
Convert the audio speech into a text transcript.
Listen to the following audio and create a corresponding text transcript.
Transform the speech into a text document.
Listen to the audio and generate a text version of it.
Create a text version based on the audio speech provided.
Table 15: Example of prompts used during training stage.
Transcribe what is said in the audio into written text. (unseen)
Extract the words from the audio and write them down. (unseen)
Produce a typed version of the audio’s spoken content. (unseen)
Capture what is being said in this audio as text. (unseen)
Hear the provided audio and provide a written text version.
Table 16: Example of prompts used during inference stage.
A.12 Efficient
The low decoding efficiency of LLMs is due to repeated invocations of the whole decoder layer during autoregressive generation. Several methods have been proposed to address this issue, including speculative sampling Leviathan et al. (2023) and KV-cache compression Zhang et al. (2023). Recently, researchers have explored using LLMs as encoders Luo et al. (2025), thereby leveraging their knowledge while improving decoding efficiency. Additionally, enhancing the computational efficiency and performance of adapters in end-to-end speech models through optimized FFN dimension design Zheng et al. (2023) represents a viable solution.

### v3

@@ 5 Discussions
Ablation study and routing method: To further validate the effectiveness of PaM, we conducted an ablation study, such as PaM without the shared expert, as detailed in Appendix A.8. We also compare PaM against a learnable routing approach without task information, commonly employed in MoE models. The results indicate that PaM is more efficient and better suited to handling multiple downstream tasks than conventional routing and fusion strategies. Moreover, Table 3 highlights the strengths of each encoder. These findings encourage researchers to select encoders tailored to their downstream tasks for the Speech LLM. While we did not incorporate this prior knowledge into the routing mechanism, leveraging it presents a promising direction for future work.
Other further work: 1.More and Unseen Tasks. Although our experiments involve three tasks and five datasets, they are representative as they encompass both semantic and acoustic-related tasks. We believe PaM can be extended to other tasks, which we will validate in future work. 2.Leverage Multimodal LLMs. Features across various encoders hinder initialization from pretrained multimodal LLMs in our work (Appendix A.9). We will explore strategies to more effectively leverage pretrained multimodal LLMs, such as Lai et al. (2024). 3.Efficiency. Improving the efficiency of LLMs has attracted much attention. Enhancing the computational efficiency of speech LLMs presents another promising avenue for exploration (Appendix A.9).
@@ 6 Related Works
Audio encoders can be classified into supervised and self-supervised models. Supervised models typically employ ASR tasks to train an end-to-end model with an audio encoder and a text decoder. By omitting the decoder, the encoder can serve as a feature extractor Radford et al. (2022); Baevski et al. (2020). Self-supervised models can be trained on unlabeled speech signals. For instance, Wav2Vec2 Baevski et al. (2020) and HuBERT Hsu et al. (2021) were trained to predict the pseudo-discrete targets at masked time steps. WavLM Chen et al. (2022) is a variant of HuBERT, designed to facilitate speaker identity extraction by using multi-speaker signals. Different model architectures, training methods, and data can result in encoders with distinct properties and advantages, making the mixture of audio encoders effective for Speech LLMs.
@@
To construct end-to-end speech LLMs, a natural approach is to extract discrete tokens from continuous speech signals and then expand the vocabulary of text LLMs to understand these speech tokens Rubenstein et al. (2023b); Veluri et al. (2024); Ma et al. (2024a). An alternative is to use an adapter layer to directly convert the continuous speech features into the continuous embedding space of the LLM. For example, QwenAudio Chu et al. (2024) employs average pooling to downsample speech features, followed by two linear layers for projection. SALMONN Tang et al. (2024) utilizes the Q-former Yu et al. (2024), a cross-attention-based adapter, to achieve a higher compression ratio. In parallel, to achieve high compression, Soundwave replaces cross-attention with a lightweight self-attention module that treats the inputs as queries Zhang et al. (2025). Compared to previous works, our adapter handles more encoders and generates different features based on the prompt, rather than a single feature for all prompts.
@@
MoE has attracted growing interest, which replaces the FFN sub-layer in Transformer models with multiple experts Shazeer et al. (2017). These MoE methods typically employ massive experts and extremely sparse activation routing, increasing model size while maintaining constant inference costs, without explicitly considering the specialization of individual experts Fedus et al. (2022); Lepikhin et al. (2020). However, the vast scale of these models presents significant challenges for deployment. In contrast, the earliest MoE research introduced a data-dependent, trainable combining method Jacobs et al. (1991); Masoudnia and Ebrahimpour (2014), which aims to decompose complex tasks into simpler sub-tasks, each managed by a dedicated expert. Such works have inspired recent advances in developing modular models called expert specialization Ma et al. (2018); Gupta et al. (2022), providing solutions for deploying large-scale MoE models Lu et al. (2024) and enabling individual experts to learn and decompose diverse knowledge Dai et al. (2024). Inspired by these insights, we propose a specialized fusion method based on MoE integrating multiple audio features to enhance Speech LLMs.
@@ 7 Conclusion
In conclusion, we propose PaM, a feature fusion method designed to provide a Speech LLM with diverse features from multiple encoders based on users’ input prompts. Experimental results indicate that PaM surpasses both single-encoder and multi-encoder baselines across a variety of tasks and datasets. We provide a detailed analysis of the feature importance of different encoders, demonstrating that PaM effectively leverages different encoders and levels of features for distinct tasks. Additionally, we present comprehensive experimental results for the selection and combination of encoders. For future work, we intend to expand the training data and incorporate additional tasks.
@@ 8 Limitations
Owing to resource constraints, our training data is limited to several hundred hours. It would be preferable to implement our method in larger-scale experiments to facilitate comparison with existing strong Speech LLMs such as Qwen-Audio Chu et al. (2024) on a more comprehensive benchmark like Air-Bench Yang et al. (2024). Additionally, we train the PaM model from scratch using a predefined list of audio encoders. It would be beneficial to investigate the addition of new encoders to an already trained Speech LLM to enhance its performance on new tasks or in new domains. We leave this for future work.
@@ A.2 Details of Models and Datasets
In this paper, we leverage multiple audio encoders and LLM to construct the end-to-end speech LLM. Our training dataset is sourced from commonly used open-source datasets, totaling approximately 450 hours of audio data, corresponding to 313,208 samples, as outlined in Table 5. For SNV, we randomly concatenate individual utterances to form new speech signals with the number of speakers ranging from one to four.
@@
Table 5: The complete training dataset.
@@
A.3 Details of Baseline Implementation
@@
For the single encoder baseline, we train the model using the same settings as in our method. For the second baseline, we train the model using the open-source codebases from WavLLM and SALMONN. We integrate the adapter components from these repositories into our code and train the baseline model using our training data, employing the same pre-trained audio encoders and LLMs as in our method. During training, we applied the same hyperparameters as our method. Since we use different encoders and LLMs compared to the baselines, we adjust the dimensions of the adapter to match the specific audio encoder and LLM we used while maintaining other dimensions independent of the audio encoders and LLM unchanged. Notably, we trained SALMONN adapter with query length=32 (as training with the original setting query length=1 failed) to ensure comparable performance with the other baseline methods.
@@ A.4 Results on Audio Caption with More Metrics
Due to the multiple evaluation metrics for the AC task, we evaluated the AC and AC QA tasks using more metrics in Table 7, including CIDEr, SPICE (with coco-caption toolkit), FENSE metric, and Sentence-BERT [^8] We used the FENSE open-source repository GitHub and scored the entire dataset with eval_system.py. For the SBERT model, we loaded the paraphrase-mpnet-base-v2 model, and for the echecker, we used echecker_clotho_audiocaps_base. However, we encountered a bug when loading the echecker model, which had an unexpected key encoder.embeddings.position_ids. To resolve this, we set strict=False. Additionally, we included results based on similarity using the SBERT model.. We found that different models exhibited similar performance trends across almost all metrics.
@@ A.5 PaM with More Encoders
We added the BEATs encoder to our framework, resulting in a total of four encoders, and found that it significantly improved the performance of our system on the AC task (Table 8). Although adding the new encoder had limited impact on the AMI and SNV tasks, incorporating the BEATs encoder improved the system’s average rank across downstream tasks (Table 9). We plan to conduct additional experiments with the EAT encoder Chen et al. (2024a) in future work.
@@ A.6 Details of Training and Inference Parameters
We train our model for five epochs with a learning rate of 5e-5, 2000 warmup steps, and bf16 precision. We freeze all encoders and the LLM, only train adapters and the fusion modules. For the LLM, we apply LoRA Hu et al. (2022) with a rank of 32 and an alpha of 64, adding LoRA only to the q_proj and k_proj. Each task has the same probability during training. During the inference stage, we select the last checkpoint on the validation set and perform greedy search.
@@ A.7 WavLLM and PaM with Larger LLM
Table 10: Results based on our PaM adapter and WavLLM with larger LLMs.
We note that the concatenation method in the Llama-based model performs significantly worse on SNV, with only 3.1% accuracy. This is because it becomes difficult to follow the instructions of the SNV task during the inference stage. After further experiments, we found that the concatenation method becomes progressively less effective on SNV. This suggests that the method struggles to achieve a balance between multitasking as training progresses.
@@ A.8 Ablation Experiments of Routing Method
  • Audio-based Routing Method. The routing method employed in most of the MoE models, such as Switch Transformers Fedus et al. (2022), DeepSeekMoE Dai et al. (2024), and MoWE Zhang et al. (2024), which use the current layer input as the routing module input, and optimize the routing module directly based on the loss of outputs. However, this approach ignores information from task labels.
@@ G $=\text{Top-}k(\text{Softmax}(\mathbf{X}))$ (11)
MoE (audio-based) 4.12 8.32 11.21 26.0% 35.66 15.61 3.17
PaM (with one expert) 4.27 10.04 13.77 45.8% 35.76 16.47 2.83
PaM (without shared) 4.31 7.89 13.43 45.2% 35.27 16.35 3.33
PaM (without task label) 3.97 7.97 13.14 43.5% 35.54 15.42 3.17
PaM (ours) 3.65 7.07 12.79 42.8% 35.47 15.70 2.50
Table 11: Ablation results on our routing method and the results based on the audio-based routing method
@@ A.9 Initialization setup for LLM
We initialize the LLM module in PaM from the open-source base model, following the setup used widely in prior work Hu et al. (2024); Ma et al. (2024b); Tang et al. (2024); Rubenstein et al. (2023a). An interesting alternative is to initialize our LLM module with the LLM module in a pretrained multimodal model like Qwen-Audio. We are not adapting such a setup because Qwen-Audio’s feature alignment was specifically designed for Whisper’s encoder, which can potentially "overfit" to features from Whisper. Our new experiments in Table 12 reveal that the feature spaces of Wav2Vec2 and WavLM are relatively similar, while Whisper’s feature space shows greater divergence. This pattern is also reflected in the weight distribution in Figure 3, where Wav2Vec2 and WavLM appear more closely aligned and significantly different from Whisper. Therefore, using other encoders and restructuring the projector module would still require re-adapting the LLM to comprehend new features.
@@ A.10 All Detailed Results
The detailed results of our experiments with multiple encoders are summarized in Table 13. We observe that, in most cases, the audio encoder that performed well on a single task also enhanced the performance of the fusion model on that task. In cases where performance degradation occurs on a specific task when using the corresponding encoder, the fusion model consistently includes the HuBERT audio encoder, suggesting that incorporating the HuBERT model may have a detrimental effect. This could be attributed to the fact that the HuBERT model is trained on a smaller pre-trained dataset compared to other audio encoders. Notably, even in this case, fusing four audio encoders yields comparable results to fusing three encoders on the AVG Rank, indicating that incorporating more encoders can still lead to performance improvements.
@@ A.11 Prompts in Training and Inference Stage
In practice, large speech-language models typically address downstream tasks using prompts that are semantically explicit but textually diverse, as illustrated in Figure 5(a). Unlike classical MoE routing methods that rely on hidden states, or other approaches such as predefined task labels (e.g., “Task ID = 3”) Kudugunta et al. (2021), and random routing Zuo et al. (2021), our model leverages natural language prompts to route inputs to the appropriate expert.
These prompts vary in phrasing but convey the same task intent. We utilize the semantic understanding capabilities of a pretrained LLM to extract task information from these prompts, rather than depending on fixed task labels. As an example, in our ASR task, we trained the routing module using 50 diverse prompts (Figure 5(b)). During evaluation, the model was tested on 200 prompts, 150 of which were unseen during training (Figure 5(c)). Our prompt-based router, powered by the LLM, achieved 100% expert selection accuracy, demonstrating strong generalization to previously unseen prompts.
Figure 5: Example of prompts.
A.12 Efficiency
The low decoding efficiency of LLMs is due to repeated invocations of the entire decoder stack during autoregressive generation. Several methods have been proposed to address this issue, including speculative sampling Leviathan et al. (2023) and KV-cache compression Zhang et al. (2023). Recently, researchers have explored using LLMs as encoders Luo et al. (2025), thereby leveraging their knowledge while improving decoding efficiency. Additionally, enhancing the computational efficiency and performance of adapters in end-to-end speech models through optimized FFN dimension design Zheng et al. (2023) represents a viable solution.